OpenStudy (anonymous):

Part 1: Create an absolute value equation that has no solutions. Explain why this equation has no solutions. Part 2: Create an absolute value equation that has one solution. Explain why this equation has one solution. Part 3: Create an absolute value equation that has two solutions. Explain why this equation has two solutions.

OpenStudy (jamesj):

In general, what do you know about |x|? What kinds of values can it have?

OpenStudy (anonymous):

idk... infinte values?

OpenStudy (jamesj):

Can it be negative, zero or positive?

OpenStudy (anonymous):

any

OpenStudy (jamesj):

No. It can never be negative.

OpenStudy (jamesj):

For example, for what value of x is |x| = -1 ?

OpenStudy (anonymous):

oohhhh

OpenStudy (jamesj):

So given that, what's an answer to your first question?

OpenStudy (anonymous):

I X I = -1 no absolute value can be negative

OpenStudy (jamesj):

yes. Now the second?

OpenStudy (anonymous):

I -9 I = 9 No matter what number it is... if its an absolute value its always positive

OpenStudy (jamesj):

that's not an equation to requires a solution. You want something like | (an expression in x) | = (a number)

OpenStudy (jamesj):

remember the definition: |x| = x, if x > 0 0, if x = 0 -x, if x < 0

OpenStudy (anonymous):

but it would still be the same right? that no matter what number it is... if its an absolute value its always positive

OpenStudy (jamesj):

yes. Let me ask you, how many solutions does |x| = 1 have?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

oohhhh... ok thats the one with 2 solutions

OpenStudy (jamesj):

So now you have an answer to part 3. What then is an answer to part 2?

OpenStudy (anonymous):

I (an expression in x) I = (a number) No matter what number it is... if its an absolute value its always positive

OpenStudy (jamesj):

|x| = -1 has no solutions |x| = 1 has two solutions |x| = (what?) has one solution.

OpenStudy (anonymous):

0

OpenStudy (jamesj):

Right. How many solutions does this equation have? |x+17| = 0

OpenStudy (anonymous):

1

OpenStudy (anonymous):

cause no matter positive or negative... it will always be 0

OpenStudy (jamesj):

Yes, good. And how many does this one have? |2x+17| = -1

OpenStudy (anonymous):

0 because its negative

OpenStudy (jamesj):

yes. For the record, what is the solution of |x+17| = 0 ?

OpenStudy (anonymous):

-17?

OpenStudy (jamesj):

yes, good.

OpenStudy (jamesj):

go back and have a look at the comments I wrote on your earlier questions. This one I think you've got under control now!